cos 36 degrees = .8090 how do I find out sin 54 degrees
Answer by jsmallt9(3759)

The key here is to recognize that 36 + 54 = 90. Or 54 = 90 -36.
So we've been given that cos(36) = 0.8090. And we're looking for sin(54) (or sin(90 - 36).

One of the properties/identities you should know is:
cos(x) = sin(90-x).
So if cos(36) = 0.8090 then sin(90-26) is also 0.8090.